# AutoWebGLM: Bootstrap And Reinforce A Large Language Model-based Web Navigating Agent

### v1

@@ Appendix A HTML Prunning Pseudo Code
Appendix B Implementation Details of AutoWebGLM
@@
Appendix C Full Results of MiniWob++
@@
Table 6: PER-TASK PERFORMANCE ON MINIWOB++
@@
Appendix D Input Prompt
Below is our input prompt for AutoWebGLM training and inference:
<html> {html_content} </html>
You are a helpful assistant that can assist with web navigation tasks.
You are given a simplified html webpage and a task description.
#Provided functions:
def click(element_id: str) -> None:
Click on the element with the specified id.
Args:
element_id: The id of the element.
def hover(element_id: str) -> None:
Hover on the element with the specified id.
Args:
element_id: The id of the element.
def select(element_id: str, option: str) -> None:
Select an option from a dropdown.
Args:
element_id: The id of the element.
option: Value of the option to select.
def type_string(element_id: str, content: str, press_enter: bool) -> None:
Type a string into the element with the specified id.
Args:
element_id: The id of the element.
content: The string to type.
press_enter: Whether to press enter after typing the string.
def scroll_page(direction: Literal[’up’, down’]) -> None:
Scroll down/up one page.
Args:
direction: The direction to scroll.
def go(direction: Literal[’forward’, backward’]) -> None:
Go forward/backward
Args:
direction: The direction to go to.
def jump_to(url: str, new_tab: bool) -> None:
Jump to the specified url.
Args:
url: The url to jump to.
new_tab: Whether to open the url in a new tab.
def switch_tab(tab_index: int) -> None:
Switch to the specified tab.
Args:
tab_index: The index of the tab to switch to.
def user_input(message: str) -> str:
Wait for user input.
Args:
message: The message to display to the user.
Returns: The user input.
def finish(answer: Optional[str]) -> None:
Finish the task (optionally with an answer).
Args:
answer: The answer to the task.
#Previous commands: {previous_commands}
#Window tabs: {exist_window_tabs_with_pointer_to_current_tab}
#Current viewport (pages): {current_position} / {max_size}
#Task: {task_description}
You should output one command to interact to the currrent webpage.
[…]
Appendix E Data Construction Prompt
Data construction prompt for web recognition description:
I want you to act as a Website Reader. Your objective is to explain a websites purpose and usage, given the websites text. Your explanation should cover all of the websites primary functions. DO NOT GUESS the purpose of the website, you SHOULD output \"None\" If you are not PRETTY sure about the purpose of the website. Note that you should only answer the purpose or usage within 20 words.
#Website Text#:
{html_content}
#Purpose#:
Data construction prompt for simple task task:
HTML:
{html_content}
Your answer SHOULD be in the following format:
Task: {Generated one-step task}
Operation: {The right operation to solve the task}
Intention: {The intention and thinking in your operation}
NOTICE:
1. Your generated task should not be too SIMPLE, NAIVE
2. You can only do \#type\# on <input> and <textarea>
Data construction prompt for complex task trace intent:
Users overall task: {task_description}
Users actions: {annotated_action_trace}
... and so on.
Appendix F Annotation Details
The annotation process was performed by 20 annotators for one month using the Google Chrome browser with our plugin installed to record their actions on assigned websites. The annotators first visited the target websites and checked whether the website descriptions matched the actual tasks. They then evaluated the tasks for clarity, relevance, achievability, complexity, and subjectivity, skipping those that didn’t meet the criteria. They carefully recorded each step during a task, including any login or captcha steps. For tasks that required an answer, the annotators manually edited the responses. If a task was not doable, they could modify its description or abandon it. We provide the demonstration of our plugin for annotation in Figure 8 and the annotation documentation in Table 7.
Figure 8: Demonstration of annotation plugin.
Table 7: Annotation Documentation.
Annotation Target We collected several websites and generated 50 unique tasks for each. For each website, annotators must perform operations according to the task descriptions to complete the requirements.
Installation of Plugin We recommend using Google Chrome for annotation.
1 Download the plugin code zip package.
2 Download the annotation data and place the folder in the directory where your plugin code is located.
3 Start Chrome, go to More > More Tools > Extensions, enable Developer mode and load the unpacked extension by selecting the unzipped plugin code directory.
Plugin Introduction To learn how to use the plugin in detail, please refer to the introduction video. The plugin has three main functions: Start/Pause Recording, End Recording, and Record Archive.
Navigation between websites and tasks is made easier with the use of ’«’ and ’»’ for previous and next websites, and ’<’ and ’>’ for previous and next tasks. The plugin also provides input fields for direct navigation.
1 Click the trash bin icon to reset the environment.
2 When encountering login or robot authentication, add an authorization record by clicking the plus (+) icon one step before the corresponding operation.
3 If the task requires an answer, click the plus (+) icon in the appropriate place to insert and edit the answer.
4 If there is no response when you type content into an input field, click the plus sign and go to the appropriate location to insert and edit the input content (note that the input field is the last object to be clicked, so make sure the last click was on an input field).
5 By default, the navigation record is marked as automatic. For manual navigation, such as typing a URL or going to the previous or next page, click the word "Navigation" in this record to change it to a manual navigation record.
6 Navigation operations are not recorded until the page is fully loaded. Wait for the navigation record to appear before you continue.
7 If you need an element on the page to hover, you can click on that element and change the "click" event to a "mouseover" record.
Annotation Process Here is a checklist for your reference.
1 Enter the assigned website number in the designated input box and navigate to the corresponding website.
2 First, verify if the auto-generated website description aligns with the website’s functionality. The description need not be detailed, but it should be generally correct. If it matches, proceed with annotation; otherwise, skip the website.
3 Task descriptions are evaluated to determine their feasibility based on criteria such as the precision of the description, whether the task is typical of the average user of the site, and its overall feasibility. Tasks that do not involve complex login or registration operations and those that do not have too many subjective conditions are preferred. If the task has minor issues, such as an unachievable condition or involving subjective conditions, you may modify it directly in the task editing box.
4 Start executing operations according to the task description, paying attention to whether there are missing or redundant operations, and handle login or robot verification as specified. After completing the task, click "End" and then "Record Archive" to save.
5 Once annotation is complete, please upload the contents of the tasks directory.
Additional Annotation Design ten tasks for each website, with each task containing at least 15 effective steps. An operation is considered effective only if it is a click, mouseover, or input change that contributes to achieving the goal. Remove any erroneous operations from the record.
1 Enter the assigned website number in the website input box and navigate to that website.
2 Briefly browse the website to determine possible tasks.
3 Choose a task number. The task box will not contain any data at this point.
4 Start recording operations.
5 After completing the operations, fill in the completed task in the task box.
6 Save the result.
Appendix G Demonstration
G.1 Weather Report
The targeted task to be executed is "What is the weather like today?". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "todays weather report"
  • Step2: Click SearchButton
  • Step3: Click SearchBar
  • Step4: Click DateButton
  • Step5: Answer
As Figure 9 shows, we end up on the webpage with a local weather report. We obtained detailed information about today’s weather as the answer, effectively completing the target task.
G.2 Shopping Advice
The targeted task to be executed is "Help me pick a Christmas gift for kids". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "Christmas gift for kids"
  • Step2: Click SearchButton
  • Step3: Click "All" tag in the category selection
  • Step4: Click SearchBar
  • Step5: Click The first product on the result page
  • Step6: Answer
As Figure 10 shows, we ultimately landed on a camera product page, where we obtained a recommendation for that camera as the answer, essentially completing the task.
G.3 Searching Article
The targeted task to be executed is "Find an article about large language model". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "large language model"
  • Step2: Click SearchButton
  • Step3: Click Wiki
  • Step4: Scroll down
  • Step5: Go backward
  • Step6: Scroll down
  • Step7: Click The link to an article
  • Step8: Answer
As Figure 11 shows, we ultimately arrived at a page featuring an article and obtained "Found a relevant article" as the answer, essentially fulfilling the task.
G.4 Searching Tools
The targeted task to be executed is "Find a tool to solve the differential equation". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "tools to solve the differential equation"
  • Step2: Click SearchButton
  • Step3: Scroll down
  • Step4: Click The link to an online tool
  • Step5: Click ODE calculator
  • Step6: Answer
As Figure 12 shows, we ultimately arrived at a page for an ODE (Ordinary Differential Equation) calculator and obtained "Found a relevant tool" as the answer, essentially completing the task.
G.5 Knowledge Query
The targeted task to be executed is "Search and tell me some basic info about the dark matter". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "dark matter"
  • Step2: Click SearchButton
  • Step3: Click Wiki
  • Step4: Scroll down
  • Step5: Answer
As Figure 13 shows, we ultimately reached a wiki page about dark matter, obtaining some basic info as the answer, and effectively completing the task.
G.6 Finding Pictures
The targeted task to be executed is "Help find a beautiful picture of the Pacific Ocean". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "Pacific Ocean Pictures"
  • Step2: Click SearchButton
  • Step3: Click A picture in the search result
  • Step4: Go backward
  • Step5: Click Another in the search result
  • Step6: Answer
As Figure 14 shows, we ultimately reached a page displaying an image of the Pacific Ocean, obtaining "Found a picture of the Pacific Ocean for you" as the answer, effectively completing the task.
G.7 Finding Research
The targeted task to be executed is "Search and tell me a hot area in AI research". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "areas in AI research"
  • Step2: Click SearchButton
  • Step3: Click A Link to a page
  • Step4: Scroll down
  • Step5: Answer
As Figure 15 shows, we ultimately reached a page about AI research, obtaining "Natural Language Processing(NLP)" as the answer, effectively completing the task.
G.8 Game Recommendation
The targeted task to be executed is "I want to play a PC game. Help me choose one". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "PC game recommendations"
  • Step2: Click SearchButton
  • Step3: Answer
As Figure 16 shows, we ultimately reached a page of searching results of games, obtaining a recommendation as the answer, effectively completing the task.
G.9 Playing Video
The targeted task to be executed is "Search and tell me a hot area in AI research". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "funny videos"
  • Step2: Click SearchButton
  • Step3: Click A Link to a vedio
  • Step4: Answer
As Figure 17 shows, we ultimately reached a page playing a funny video, effectively completing the task.
G.10 Online Shopping Assistance with Pop-Up Interruption
The targeted task to be executed is "Find and select a highly rated toaster". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "best toaster 2024"
  • Step2: Click SearchButton
  • Step3: Click a link from the search results leading to an online shopping site (Encounter a pop-up asking to subscribe to the newsletter.)
  • Step4: Scroll down, but the interaction is blocked by the pop-up
  • Step5: Answer
As Figure 18 shows, we do not reach the intended product selection. The presence of an unexpected pop-up interrupts the task execution, demonstrating the system’s limitation in handling unexpected graphical elements and pop-ups. This outcome underscores the need for enhanced capabilities in graphical recognition and interaction handling, ensuring smoother navigation and task completion on web pages with complex elements.
G.11 Knowledge Query with Hallucination
The targeted task to be executed is "Tell me some basic info about NLP". The actual execution steps can be summarized as follows:
  • Step1: Answer
As Figure 19 shows, this case is a classic example of the hallucination fallacy, where the system responded directly without going through the webpage, and the response came from the hallucination rather than the webpage information
G.12 Technological Breakthrough Summary with misinterpretation
The targeted task to be executed is "Summarize a recent technological breakthrough in renewable energy". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "latest technological breakthrough in renewable energy 2024"
  • Step2: Click SearchButton
  • Step3: Click a link in search results (The system selects a link to a general overview of renewable energy technologies instead of a specific article on recent breakthroughs.)
  • Step4: Scroll down
  • Step5: Answer
As Figure 20 shows, we do not reach the intended outcome. Instead of summarizing a recent technological breakthrough, the system provides a generalized overview of renewable energy. This outcome highlights a misinterpretation of task context, demonstrating the system’s challenge in distinguishing between general information and specific, recent developments.
G.13 Map Query with Poor Graphical Recognition
The targeted task to be executed is "Where is Beijing relative to Shanghai according to the map". The actual execution steps can be summarized as follows:
  • Step1: Type SearchBar "Beijing"
  • Step2: Click SearchButton
  • Step3: Answer
As Figure 21 shows, we ultimately reached a page displaying a description of Beijing and the Beijing map, obtaining "Northside" as the answer. The answer is to some extent too simple. This case illustrates two minor flaws in our system: one is that it has a slight lack of understanding of the graphical interface, and the other is that it sometimes hallucinates, and the answers it gets are not always from web information.
Figure 9: Weather Report
Figure 10: Shopping Advice
Figure 11: Searching Article
Figure 12: Searching tools
Figure 13: Knowledge Query
Figure 14: Finding Pictures
Figure 15: Finding Research
Figure 16: Game Recommendation
Figure 17: Playing Video
Figure 18: Online Shopping Assistance with Pop-Up Interruption
Figure 19: Knowledge Query with Hallucination
Figure 20: Technological Breakthrough Summary with misinterpretation
Figure 21: Map Query with Poor Graphical Recognition

### v2

@@ 3.2.1. Observation space
HTML. The HTML webpages are vast and complex, so it is necessary to simplify them before inputting them into the model. The simplification process aims to extract essential information while eliminating redundant or disruptive elements that could hinder the model’s understanding. Throughout this process, the HTML’s basic structure and significant content information must be retained to enable the model to comprehend and utilize this information for effective web browsing. HTML Pruner can efficiently convert a tree of elements into a concise representation. We can use the processing techniques to streamline the original HTML format into a more understandable structure for the model to interpret and manage, improving model effectiveness in web browsing tasks.
3.2.2. Action space
As the approach of this work is to build a language model-based web browsing agent, we focus on operational possibilities when constructing the action space. On an extensive summary of experiences in the real task execution process, we define a complete and self-consistent action space (in Table 1) for the language model to act in the web browsing world. We design our prompt input in Section B.
4. Building AutoWebGLM
Figure 4. Data Construction. Data construction is divided into two main stages; the first stage is webpage recognition tasks and simple tasks operation construction, and the second stage is complex tasks construction.
In this section, we detail the construction of a web browsing agent. Given the high costs associated with manual data construction and the inadequacy of current LLMs for automated data generation, we employed a Human-AI hybrid data construction method to efficiently produce large volumes of training data at a reduced cost. Additionally, we implemented a multi-stage training approach, rather than relying solely on imitation learning, to enhance our model’s general and specialized web browsing capabilities.
4.1. Data Construction
As illustrated in Figure 4, we suggest a hybrid human-AI Data Construction method to create our training data in response to these challenges. After careful consideration, we categorize our data into two types for construction:
4.1.1. Web Recognition & Simple Task Operation Construction
4.1.2. Complex Task Operation Construction
We first designed 50 complex tasks for each website using the prompting technique referring to Evol-Instruct (Xu et al., 2023b), from which about 20 feasible tasks were manually selected and labeled. For operation sequence, due to the high complexity of the tasks, even the most advanced LLMs cannot complete the task with satisfactory accuracy. Therefore, we leveraged manual annotations to capture web task executions via a browser plugin that records actions during website tasks. Chain-of-thought (Wei et al., 2022) reasoning has been proven to improve task comprehension and model performance (Kojima et al., 2022; Wang et al., 2022) significantly. However, leveraging human annotators to document their intent and reasoning during web browsing is inefficient. To improve the CoT construction process, we used GPT-4 as the operational intent predictor. Our first approach of iterative step-by-step creation proved to generate weak operational links and incurred high API costs due to data construction. To address this, we employed a global thought chain prompting method, where all operations and critical HTML segments are inputted into a trace. Then, we prompted GPT-4 to output intentions for each step. This method improves the accuracy and cohesion of each step, thus forming highly relevant, consistent thought chains.
Figure 5. Dataset Proportion. Piechart of the distribution of splits within our training data.
Figure 6. The Training Procedure. First, the model learns webpage interpretation and operation via curriculum learning. Next, it self-samples training data, learning from its mistakes. Finally, it self-plays in the environment, becoming a domain expert.
4.2. Training
We train the model through three steps illustrated in Figure 6.
4.2.1. Step 1: Curriculum Learning
(1) $\mathcal{L_{SFT}}(\pi_{\theta})=-\mathbb{E}_{(x,y)\thicksim\mathcal{D}}\left[\log\pi_{\theta}(y\mid x)\right]$
This approach enhances the model’s comprehension of webpages and its capability as an agent to perform operations within the environments. Significantly, we use curriculum learning (CL), which mimics the human learning process, advocating for models to start learning from easy samples and gradually advance to complex ones. It has been demonstrated in prior works(Bengio et al., 2009; Wang et al., 2021) to improve model capabilities substantially.
4.2.2. Step 2: Reinforcement Learning
After constructing contrastive data $D_{\text{Const.}}$, we employ the DPO(Rafailov et al., 2023) training approach to make $M_{\text{SFT}}$ learn from its mistakes and further enhance its capabilities. During the training, we found that the direct use of DPO loss led to instability. To mitigate this issue, we propose including SFT loss to stabilize the reinforcement learning process and increase the number of training steps while ensuring no loss of the original model’s natural language and agent abilities, achieving a more robust model $M_{\text{DPO}}$:
(2) $\begin{array}[]{l}\mathcal{L}_{DPO}(\pi_{\theta};\pi_{\mathrm{ref}})=\\ -\mathbb{E}_{(x,y_{w},y_{l})\thicksim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}\right.\right.\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right]\end{array}$
(3) $\mathcal{L}_{SFT}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\thicksim\mathcal{D}}\left[\log\pi_{\theta}(y_{w}\mid x)\right]$
(4) $\mathcal{L}_{Total}=\lambda\cdot\mathcal{L}_{DPO}+\mathcal{L}_{SFT}$
4.2.3. Step 3: Rejection Sampling Finetuning
4.3. Benchmark: AutoWebBench
5. Experiments
5.1. Main Results
AutoWebBench. As discussed in Section 4.3, We divide the test set into four splits: Chinese, English, in-domain, and out-of-domain, for evaluation purposes. We use the Step Success Rate (SSR) as our evaluation metric. The results are in Table 2.
Table 2. The performance on AutoWebBench.
English Chinese
Model Size Cross- Task Cross- Domain Cross- Task Cross- Domain
GPT-3.5-Turbo N/A 12.1 6.4 13.5 10.8
GPT-4 N/A 38.6 39.7 36.7 36.3
Claude2 N/A 13.2 8.1 13.0 7.9
LLaMA2 7B 3.3 2.5 - -
LLaMA2 70B 8.3 8.9 - -
Qwen 7B 9.0 7.6 9.1 7.5
AutoWebGLM 6B 64.8 58.6 65.4 61.8
Mind2Web. We use the settings from Mind2Web with SSR as our primary evaluation metric. To compare the model fairly, we utilize the MindAct framework provided by Mind2Web to evaluate the model’s performance. The results are in Table 3.
Table 3. The performance on Mind2Web. † indicates that only top-10 candidates were used for this test, otherwise top-50 was used. * indicates model’s finetuning on train set.
Model Size Cross- Task Cross- Website Cross- Domain Average
GPT-3.5-Turbo N/A 17.4 16.2 18.6 17.4
GPT-4 N/A 36.2 30.1 26.4 30.9
Flan-T5-XL* 3B 52.0 38.9 39.6 43.5
Html-T5-XL* 543B 71.5 62.2 67.1 66.9
LLaMA2* 7B 52.7 47.1 50.3 50.1
LLaMA2* 70B 55.8 51.6 55.7 54.4
Qwen-VL* 9.6B 12.6 10.1 8.0 10.2
SeeClick* 9.6B 23.7 18.8 20.2 20.9
AutoWebGLM 6B 66.4 56.4 55.8 59.5
MiniWoB++ & WebArena. For MiniWob++, following the experimental setup from WebAgent (Gur et al., 2023), we test MiniWoB++ with 56 tasks by running 100 evaluation episodes per task to evaluate model capabilities. For WebArena, we integrate our HTML parser module and action execution module into the WebArena environment to make it compatible with our system. The results are in Table 4.
Table 4. The performance on MiniWoB++ and WebArena. * indicates model’s finetuning on train set.
Model Size MiniWoB++ WebArena
GPT-3.5-Turbo N/A 13.4 6.2
GPT-4 N/A 32.1 14.4
Text-Bison-001 N/A - 5.1
LLaMA2 7B 42.8* 1.2
LLaMA2 70B 47.1* 0.6
Html-T5-XL 543B 85.6* -
WebN-T5-XL 3B 48.4* -
Lemur 70B - 5.3
AutoWebGLM 6B 89.3 18.2
5.2. System Execution Efficiency
Table 5. System Execution Efficiency
Action count/tr Fetch Parse Predict Execute Loading
type_string 1.00 362.00 28.88 2282.99 6.12 2082.61
click 3.62 438.62 72.07 2252.10 28.10 2094.60
finish 0.38 419.33 66.00 3054.16 6.00 2119.06
scroll_page 0.38 475.33 88.67 3396.37 5.33 2033.35
Others 0.25 680.50 152.50 2666.08 10.00 2142.92
Average - 436.93 68.68 2407.34 20.36 2092.13
Percentage(%) - 8.70 1.37 47.89 0.41 41.64
Furthermore, since execution speed is critical to user experience, we conduct a series of performance experiments to evaluate the execution efficiency of each system component and identify areas that could be further optimized. The results of these experiments are presented in Table 5.
5.3. Ablation Study
Table 6. Ablation study. AutoWebBench and WebArena do not have a training set, while the RFT stage is only suitable for sampling in the environment, so we represent them by ”-”.
Method AutoWebBench Mind2Web MiniWob++ WebArena
Training Data Ablation
Only Train Set - 48.1 44.3 -
+) Stage1 23.5 48.4 48.3 2.5
+) Stage2 60.2 55.2 78.9 7.6
+) Stage1+2 61.8 56.7 81.7 8.3
Training Strategy Ablation
SFT 61.8 56.7 81.7 8.3
+) DPO 62.7 59.5 80.8 8.5
+) RFT - - 89.3 18.2
AutoWebGLM 62.7 59.5 89.3 18.2
5.4. Case Study and Error Analysis
To assess the effectiveness of our model, we conduct a series of case studies covering a range of web-based tasks, including everyday use, leisure and relaxation, and academic research, covering the typical range of web requirements. Our system achieves satisfactory results in most scenarios.
Table 7. Error Distribution in Web Task Automation
Error Type Proportion
Hallucinations 44%
Poor Graphical Recognition 28%
Misinterpretation of Task Context 20%
Pop-Up Interruption 8%
6. Future Direction
6.1. Multimodal Input
While HTML input has produced satisfactory results in many scenarios, our system falters when confronted with advanced web applications such as maps, animations, and video browsing. In our analysis, the strength of image input lies in its indispensable role in interpreting images, icons, and special effects. However, compared to text input, image input presents additional challenges in understanding numerals and extensive web text. Consequently, we consider that a multimodal system, integrating HTML and webpage screenshots, combines the advantages of both modalities, substantially enhancing the model’s capability in web browsing tasks.
6.2. Reasoning and Self-check Techniques
The system’s efficiency and success rate in web browsing may decrease when dealing with unfamiliar websites or those with unique operating logic. To mitigate this issue, an exciting avenue for exploration is the development of novel reasoning strategies distinct from the Chain-of-Thought approach, enabling the model to make better-informed decisions based on previous browsing experiences, thereby improving the success rate and efficiency of web browsing. Moreover, due to unstable internet connections and other factors, the stability of a real web environment is not guaranteed. Thus, self-check mechanisms within the web browsing agent system, including confirming the current state and verifying the intended operation’s effect, could significantly improve the system’s robustness and effectiveness.
6.3. Mobile Application
The mobile platform is another promising application scenario with massive potential. Compared to the web platform, it presents its challenges and opportunities. For example, due to their screen size, mobile devices display fewer elements within the viewport, simplifying the page XML. Furthermore, the operation logic on mobile platforms is generally more straightforward than on web platforms. However, mobile operation space includes more complex actions such as gestures, and mobile platforms face more system security restrictions, imposing additional constraints on software development.
7. Conclusion
In this work, we present AutoWebGLM, an advanced language model-based agent exhibiting robust performance in various autonomous web navigation benchmarks. Our model addresses extant LLM limitations and simplifies webpages by effectively controlling HTML text length and handling the web’s open-domain nature. We strategically employ curriculum learning, reinforcement learning, and rejection sampling finetuning to enhance webpage comprehension and browser operation learning. We also introduce a unique bilingual web browsing benchmark— that lays a solid foundation for future research. Our findings represent significant progress in utilizing LLMs for intelligent agent tasks.
Acknowledgements.
This work is supported by Natural Science Foundation of China (NSFC) 62276148 and 62425601, the New Cornerstone Science Foundation through the XPLORER PRIZE and Tsinghua University (Department of Computer Science and Technology) -Siemens Ltd., China Joint Research Center for Industrial Intelligence and Internet of Things (JCIIOT).
Appendix A Implementation Details of AutoWebGLM
@@
Appendix B Input Prompt
Below is our input prompt for AutoWebGLM:
<html> {html_content} </html>
You are a helpful assistant that can assist with web navigation tasks.
You are given a simplified html webpage and a task description.
#Provided functions:
def click(element_id: str) -> None:
Click on the element with the specified id.
Args:
element_id: The id of the element.
... (Other function definitions)
#Previous commands: {previous_commands}
#Window tabs: {exist_window_tabs_with_pointer_to_current_tab}
#Current viewport (pages): {current_position} / {max_size}
#Task: {task_description}
You should output one command to interact to the currrent webpage.
[…]
Appendix C Data Construction Prompt
Data construction prompt for task and trace intent:
HTML:
{html_content}
Your answer SHOULD be in the following format:
Task: {Generated one-step task}
Operation: {The right operation to solve the task}
Intention: {The intention and thinking in your operation}
NOTICE:
1. Your generated task should not be too SIMPLE, NAIVE
2. You can only do \#type\# on <input> and <textarea>
Users overall task: {task_description}
Users actions: {annotated_action_trace}
... and so on.
Appendix D Annotation Details
The annotation process was performed by 20 annotators for one month using the Google Chrome browser with our plugin installed to record their actions on assigned websites. The annotators first visited the target websites and checked whether the website descriptions matched the actual tasks. They then evaluated the tasks for clarity, relevance, achievability, complexity, and subjectivity, skipping those that didn’t meet the criteria. They carefully recorded each step during a task, including any login or captcha steps. For tasks that required an answer, the annotators manually edited the responses. If a task was not doable, they could modify its description or abandon it.
Appendix E Full Results of MiniWob++
@@
Table 8. PER-TASK PERFORMANCE ON MINIWOB++